# Lazy-DaSH: Lazy Approach for Hypergraph-based Multi-robot Task and Motion Planning

### v1

@@ IV The Lazy-DaSH Method
The planning process of Lazy-DaSH starts by entering the task planning layer, first constructing the task space representation, or task space hypergraph ($\mathcal{H}_{\mathcal{T}}$), and subsequently generating a discrete task plan represented as task-extended hypergraph ($\mathcal{H}_{\texttt{TE}}$). The resulting task plan is an unvalidated schedule that presumes all motions within the task space elements are feasible and lazily validates motion feasibility in the subsequent motion planning layer. The motion planning layer first constructs the motion representation, or motion hypergraph ($\mathcal{H}_{\mathcal{M}}$), only for the task space elements identified as relevant by the task plan (unvalidated schedule). The motion feasibility is evaluated during the motion query phase by generating the motion-extended hypergraph ($\mathcal{H}_{\texttt{ME}}$). This approach alleviates the need for an unnecessarily expensive motion feasibility check for actions not included in the task plan or occurring after an action with no valid motion. The resulting motion plan is an optimistic schedule, requiring the conflict resolution layer to generate the collision-free valid schedule.
@@
1:Initial task space element set $T_{\texttt{init}}$, goal task space element set $T_{\texttt{goal}}$, valid transition set $A$
@@
3:$\mathcal{S}_{\mathcal{T}},\mathcal{S}_{\mathcal{M}}\leftarrow\emptyset$triangleright task and motion schedules
4:$\mathcal{C}_{\mathcal{T}},\mathcal{C}_{\mathcal{M}}\leftarrow\emptyset$$\triangleright$ task and motion constraints
@@
7:    $\mathcal{H}_{\mathcal{T}}\leftarrow\textsc{TaskSpaceHg}(\mathcal{H}_{\mathcal{T}},A)$
@@
10:    while $\mathcal{S}_{\mathcal{T}}$ valid and $S_{\mathcal{M}}$ not valid
11:          // Generate Motion Hypergraph
12:         $\mathcal{H}_{\mathcal{M}},\mathcal{C}_{\mathcal{T}},\mathcal{C}_{\mathcal{M}}\leftarrow\textsc{MotionHg}(\mathcal{S}_{\mathcal{T}},\mathcal{H}_{\mathcal{T}},\mathcal{H}_{\mathcal{M}},\mathcal{C}_{\mathcal{M}})$
13:          // Query Optimistic Schedule
14:         $\mathcal{S}_{\mathcal{M}},\mathcal{C}_{\mathcal{T}}\leftarrow\textsc{QueryMotionPlan}(\mathcal{S}_{\mathcal{T}},\mathcal{H}_{\mathcal{T}},\mathcal{H}_{\mathcal{M}},\mathcal{C}_{\mathcal{M}})$
15:         if $\mathcal{C}_{\mathcal{T}}\neq\emptyset$
16:             break
17:          // Generate Valid Schedule
18:         $\mathcal{S}_{\mathcal{M}},\mathcal{C}_{\mathcal{M}}\leftarrow\textsc{ResolveConflicts}(\mathcal{S}_{\mathcal{M}},\mathcal{C}_{\mathcal{M}})$
19:return $\mathcal{S}_{\mathcal{M}}$
IV-A Task Planning Layer
The highest layer of the hierarchy, the task planning layer, encompasses the construction of the task representation and its query phases within the task-level domain.
IV-A1 Representation Construction (Task Space Hypergraph)
@@
[…]
Generating $\mathcal{H}_{\mathcal{T}}=(\mathcal{V}_{\mathcal{T}},\mathcal{E}_{\mathcal{T}})$ begins by connecting the set of allowed task space elements (robots, objects with locations, and robots with objects) with all possible transitions. The objects with location task space elements include the start (and goal vertex) for each object, which are all connected with a single hyperarc to task source vertex $v_{\mathcal{T}}^{\texttt{source}}$ (and task sink vertex $v_{\mathcal{T}}^{\texttt{sink}}$), to capture the full scope of the start (and goal) task states. The reachability of objects in the start and goal locations is then evaluated based on the robot’s capabilities (i.e., maximum payload and range) or grasp pose (i.e., a valid solution of inverse kinematics). However, the reachability only indicates that the object is within the graspable range of some robots, without considering the motion feasibility of the robot reaching the object. Feasibility will be evaluated at the motion planning layer by running a motion planner.
@@
IV-A2 Query (Task-extended Hypergraph)
@@
Since the hyperarcs in $\mathcal{H}_{\mathcal{T}}$ transition between different task space compositions, directly querying the task plan within $\mathcal{H}_{\mathcal{T}}$ may result in movable entities appearing in multiple task space compositions simultaneously. This may cause a task space element to perform multiple transitions at the same time, leading to an infeasible plan. To avoid this issue, we generate the task-extended hypergraph $\mathcal{H}_{\texttt{TE}}$, sequentially expanding the hyperarcs from the start task vertices by selectively choosing a viable transition for the frontiers in the current transition history.
@@
The task-extended hypergraph is defined as $\mathcal{H}_{\texttt{TE}}=(\mathcal{V}_{\texttt{TE}},\mathcal{E}_{\texttt{TE}})$. Each task-extended vertex $v_{\texttt{{TE}}}=\langle v_{\mathcal{T}},\Pi_{v_{\texttt{{TE}}}^{\texttt{{source}}}v_{\texttt{{TE}}}}\rangle$ $\in\mathcal{V}_{\texttt{TE}}$ is defined by a task vertex $v_{\mathcal{T}}\in\mathcal{V}_{\mathcal{T}}$ and a task-extended transition history $\Pi_{v_{\texttt{{TE}}}^{\texttt{{source}}}v_{\texttt{{TE}}}}$ that stores the history connecting from task-extended source vertex $v_{\texttt{{TE}}}^{\texttt{{source}}}=\langle v_{\mathcal{T}}^{\texttt{source}},\emptyset\rangle$ to $v_{\texttt{{TE}}}$. Each task-extended hyperarc $E_{\texttt{TE}}=\langle\texttt{Tail},\texttt{Head},E_{\mathcal{T}}\rangle\in\mathcal{E}_{\texttt{TE}}$ includes the information about the tail, head, and task hyperarc that contributes to the history transitions. Beginning from $v_{\texttt{TE}}^{\texttt{source}}$, the search process finishes when the task-extended hyperarc finds the task-extended sink vertex $v_{\texttt{TE}}^{\texttt{sink}}=\langle v_{\mathcal{T}}^{\texttt{sink}},\Pi_{v_{\texttt{TE}}^{\texttt{{source}}}v_{\texttt{TE}}^{\texttt{{sink}}}}\rangle$ in a head set. The task plan is obtained by extracting the task vertices $v_{\texttt{TE}}.v_{\mathcal{T}}$ from each $v_{\texttt{TE}}$ along $\Pi_{v_{\texttt{TE}}^{\texttt{{source}}}v_{\texttt{TE}}^{\texttt{{sink}}}}$. The resulting task plan is an unvalidated schedule, which does not account for the motion feasibility along the task transitions.
@@
IV-B Motion Planning Layer
@@
This section describes the second layer of the hierarchy, the motion planning layer, which involves the construction of the motion representation and the query phases within the motion-level domain.
IV-B1 Representation Construction (Motion Hypergraph)
The motion hypergraph $\mathcal{H}_{\mathcal{M}}=(\mathcal{V}_{\mathcal{M}},\mathcal{E}_{\mathcal{M}})$ captures the motion details of $\mathcal{H}_{\mathcal{T}}$. Each vertex $v_{\mathcal{M}}=\langle v_{\mathcal{T}},q\rangle\in\mathcal{V}_{\mathcal{M}}$ contains both a task space vertex $v_{\mathcal{T}}$ and an explicit configuration $q$, and each motion hyperarc $E_{\mathcal{M}}=\langle\texttt{Tail},\texttt{Head},E_{\mathcal{T}},\pi\rangle\in\mathcal{E}_{\mathcal{M}}$ consists of its tail and head sets, a corresponding task hyperarc $E_{\mathcal{T}}$, and a configuration path $\pi$ connecting motion vertices $E_{\mathcal{M}}.\texttt{Tail}$ and $E_{\mathcal{M}}.\texttt{Head}$. As in DaSH [2], the motion hyperarcs can be categorized as a composite $E_{\mathcal{M}}^{\texttt{comp}}$, transition $E_{\mathcal{M}}^{\texttt{tran}}$, or move hyperarc $E_{\mathcal{M}}^{\texttt{move}}$. In the object rearrangement problem, for instance, a motion for merging or diverging of task space elements due to grasping or hand-over is a motion transition hyperarc $E_{\mathcal{M}}^{\texttt{trans}}$, and a motion of the manipulator or manipulator holding an object within the same task space elements is a motion move hyperarc $E_{\mathcal{M}}^{\texttt{move}}$.
The feasibility of each $E_{\mathcal{M}}$ reflects the feasibility of $E_{\mathcal{M}}.\pi$ and is validated in a lazy manner by querying a lazy sampling-based motion planner such as Lazy-PRM [5]. The edge-unvalidated roadmap generated by lazy sampling-based motion planners is validated in the order specified by the unvalidated schedule. If a motion query fails while tracing the unvalidated schedule, either a motion constraint is introduced or additional vertices are sampled to find an alternative path in the roadmap. The lazy motion validation approach minimizes unnecessary computational effort by deferring validation until it becomes necessary. If an invalid motion is found in the unvalidated schedule, it implies that the validity of the subsequent motions does not need to be checked yet as the task plan may change.
The construction of $\mathcal{H}_{\mathcal{M}}$ is similar to the motion hypergraph construction stage introduced in Section IV-C of [2], but our approach features a lazy motion validation during the construction. Initially, the transition sampler determines transition configurations (e.g., grasp and hand-over pose) only for transitions between the relevant task space elements. Then edge-unvalidated roadmaps are generated within the relevant task space elements to plan motion for $E_{\mathcal{M}}^{\texttt{trans}}$ and $E_{\mathcal{M}}^{\texttt{move}}$. If the transition sampler fails to generate valid transition configurations, the construction process stops and creates a task constraint to exclude the infeasible interaction during the replanning of the task plan.
IV-B2 Query (Motion-extended Hypergraph)
Similar to $\mathcal{H}_{\texttt{TE}}$, the motion sequence is queried by the motion-extended hypergraph $\mathcal{H}_{\texttt{ME}}$ to avoid movable entities attempting multiple motions simultaneously. Since the abstract guideline of the motion plan is defined by the task plan, the motion-extended hypergraph is basically a motion-detailed version of the task plan. The feasibility of the motions between these transition configurations is validated while tracing the task plan. If motion validation fails after a certain amount of effort (e.g., timeout), the task constraints are provided to the task query phase to enforce restrictions on the unsuccessful task sequence, prompting a replanning of the task plan.

### v2

@@ IV-A Overview
The planning process of Lazy-DaSH starts by entering the task planning layer, first constructing the task space representation, or task space hypergraph ($\mathcal{H}_{\mathcal{T}}$), and subsequently generating a discrete task plan represented as task-extended hypergraph ($\mathcal{H}_{\texttt{TE}}$). The resulting task plan is an unvalidated schedule that presumes all motions within the task space elements are feasible. The unvalidated schedule is first checked in the task conflict detection layer against the static object poses and corresponding robot configurations in the unvalidated history. Any invalid task sequences are then corrected by replanning with task constraints. If a task query fails due to conflicts in task constraints, the task space representation is expanded to broaden the search space. This approach reduces subsequent motion planning calls and collision checks in the conflict resolution layer by first identifying infeasible task sequences in advance. The feasibility of motions within the task plan is then lazily validated in the motion planning layer. The motion planning layer first constructs the motion representation, or motion hypergraph ($\mathcal{H}_{\mathcal{M}}$), only for the elements of the task space identified as relevant by the task plan (unvalidated schedule). The motion feasibility is evaluated during the motion query phase by generating the motion-extended hypergraph ($\mathcal{H}_{\texttt{ME}}$). This approach alleviates the need for an unnecessarily expensive motion feasibility check for actions not included in the task plan or occurring after an action with no valid motion. The resulting motion plan is an optimistic schedule, requiring the motion conflict resolution layer to generate the collision-free valid schedule.
@@
1:Initial task space element set $T_{\texttt{init}}$, goal task space element set $T_{\texttt{goal}}$, valid transition set AA
@@
[…]
@@
7:  $\mathcal{H}_{\mathcal{T}}\leftarrow\textsc{TaskSpaceHg}(\mathcal{H}_{\mathcal{T}},\mathcal{C}_{\mathcal{T}},A)$
@@
10:   // Task-Vertex Conflict Check
11:  $\mathcal{C}_{\mathcal{T}}\leftarrow\textsc{DetectTaskConflicts}(\mathcal{S}_{\mathcal{T}})$
12:  while $\mathcal{S}_{\mathcal{T}}$ valid and $S_{\mathcal{M}}$ not valid
13:     // Generate Motion Hypergraph
14:    $\mathcal{H}_{\mathcal{M}},\mathcal{C}_{\mathcal{T}},\mathcal{C}_{\mathcal{M}}\leftarrow\textsc{MotionHg}(\mathcal{S}_{\mathcal{T}},\mathcal{H}_{\mathcal{T}},\mathcal{H}_{\mathcal{M}},\mathcal{C}_{\mathcal{M}})$
15:     // Query Optimistic Schedule
16:    $\mathcal{S}_{\mathcal{M}},\mathcal{C}_{\mathcal{T}}\leftarrow\textsc{QueryMotionPlan}(\mathcal{S}_{\mathcal{T}},\mathcal{H}_{\mathcal{T}},\mathcal{H}_{\mathcal{M}},\mathcal{C}_{\mathcal{M}})$
17:    if $\mathcal{C}_{\mathcal{T}}\neq\emptyset$
18:     break
19:     // Generate Valid Schedule
20:    $\mathcal{S}_{\mathcal{M}},\mathcal{C}_{\mathcal{M}}\leftarrow\textsc{MotionConflicts}(\mathcal{S}_{\mathcal{M}},\mathcal{C}_{\mathcal{M}})$
21:return $\mathcal{S}_{\mathcal{M}}$
IV-B Task Planning Layer
The highest layer of the hierarchy, the task planning layer, encompasses the construction of the task representation and its query phases within the task-level domain. In an effort to keep the search space small, the initial set of object states only includes the start and goal position for each object. This is a greedy assumption that no intermediate object positions are required (which is the most common case). When that assumption fails, the method samples new object positions which satisfy the stability constraints and extends the representation to encode these object states.
IV-B1 Representation Construction (Task Space Hypergraph)
@@
For the multi-manipulator object rearrangement problems considered in this paper, the task vertices in $\mathcal{H}_{\mathcal{T}}$ representing the task space elements correspond to robots by themselves ($R_{1}$ and $R_{2}$ in Fig. 2), robots holding objects ($R_{1}\!+\!O_{1}$ and $R_{2}\!+\!O_{1}$), or objects at a particular location as indicated by the constraints in the task space element ($O_{1}^{s}$ and $O_{1}^{g}$). In Lazy-DaSH, the object-only task space elements differ from those in DaSH, where objects are associated with a more “location-specific” at configuration $q$ under constraints $s_{c}(O,q)$. The location-specific constraints used here allow object-only task space elements to encode explicit object locations at the task space level ($O_{1}^{s}$ and $O_{1}^{g}$ in Fig. 2, representing the start and goal, respectively). This formulation enables $\mathcal{H}_{\mathcal{T}}$ to be queried for generating a task plan (unvalidated schedule) that links start and goal object-only task space elements.
As in DaSH [2], generating $\mathcal{H}_{\mathcal{T}}=(\mathcal{V}_{\mathcal{T}},\mathcal{E}_{\mathcal{T}})$ begins with the initial task space elements (robot-only and object-only with stable poses) and recursively applies predefined transitions such as pick, place, and handover. Each transition couples or decouples task space elements according to their constraints, and the expansion continues until no additional feasible task space elements can be generated. The initial object task space elements include the start (and goal vertex) for each object, which are all connected with a single hyperarc to task source vertex $v_{\mathcal{T}}^{\texttt{src}}$ (and task sink vertex $v_{\mathcal{T}}^{\texttt{sink}}$), to capture the full scope of the start (and goal) task states. The reachability of objects in the start and goal locations is then evaluated based on the robot’s capabilities (i.e., maximum payload and range) or grasp pose (i.e., a valid solution of inverse kinematics). The grasp pose may be evaluated across multiple candidates, such as sides, top, or other feasible configurations. However, the reachability only indicates that the object is within the graspable range of some robots, without considering the motion feasibility of the robot reaching the object. Feasibility will be evaluated at the motion planning layer by running a motion planner. When the task space expansion extension is requested from the subsequent layer, specifically the task conflict detection layer, object-only task space elements are generated, and the $\mathcal{H}_{\mathcal{T}}$ is reconstructed to include these new elements while ensuring valid transition rules (e.g., grasp and handover) are satisfied.
@@
IV-B2 Query (Task-extended Hypergraph)
@@
Since the hyperarcs in $\mathcal{H}_{\mathcal{T}}$ transition between different task space compositions, directly querying the task plan within $\mathcal{H}_{\mathcal{T}}$ may result in movable entities appearing in multiple task space compositions simultaneously. This may cause a task space element to perform multiple transitions at the same time, leading to an infeasible plan. To avoid this issue, we maintain a set of partial task hyperarcs, which represent potential transitions that are candidates for expansion but remain unexpandable until all of their tail task space elements are satisfied in the frontier of the current transition history. A hyperarc is expanded only if it does not result in a vertex with multiple outgoing hyperarcs. In DaSH [2], which performs integrated task and motion query, these partial hyperarcs are maintained within the motion hyperarcs, which enumerate more exhaustive options, including motions confined to the same task space elements as well as transitions for interacting with other robots and objects. In contrast, Lazy-DaSH preserves partial hyperarcs by abstracting motions and emphasizing task-oriented transitions. The expansion process is recorded by generating the task-extended hypergraph $\mathcal{H}_{\texttt{TE}}$, where hyperarcs are sequentially expanded from the start task vertices by selectively choosing viable transitions from the current transition history. This enables independent task threads to be explored in parallel, supporting asynchronous execution.
This method offers a distinct advantage in handling task constraints. If a potential transition is found to be infeasible, its corresponding partial hyperarc can be pruned without affecting other valid, parallel planning threads. Such flexibility is difficult to achieve in coupled planning approaches, which construct a single sequential plan and cannot easily modify independent branches of the search.
@@
The task-extended hypergraph is defined as $\mathcal{H}_{\texttt{TE}}=(\mathcal{V}_{\texttt{TE}},\mathcal{E}_{\texttt{TE}})$. Each task-extended vertex $v_{\texttt{{TE}}}=\langle v_{\mathcal{T}},\Pi_{v_{\texttt{{TE}}}^{\texttt{{src}}}v_{\texttt{{TE}}}}\rangle$ $\in\mathcal{V}_{\texttt{TE}}$ is defined by a task vertex $v_{\mathcal{T}}\in\mathcal{V}_{\mathcal{T}}$ and a task-extended transition history $\Pi_{v_{\texttt{{TE}}}^{\texttt{{src}}}v_{\texttt{{TE}}}}$ that stores the history connecting from task-extended source vertex $v_{\texttt{{TE}}}^{\texttt{{src}}}=\langle v_{\mathcal{T}}^{\texttt{src}},\emptyset\rangle$ to $v_{\texttt{{TE}}}$. Each task-extended hyperarc $E_{\texttt{TE}}=\langle\texttt{Tail},\texttt{Head},E_{\mathcal{T}}\rangle\in\mathcal{E}_{\texttt{TE}}$ includes the information about the tail, head, and task hyperarc that contributes to the history transitions. Beginning from $v_{\texttt{TE}}^{\texttt{src}}$, the search process finishes when the task-extended hyperarc finds the task-extended sink vertex $v_{\texttt{TE}}^{\texttt{sink}}=\langle v_{\mathcal{T}}^{\texttt{sink}},\Pi_{v_{\texttt{TE}}^{\texttt{{src}}}v_{\texttt{TE}}^{\texttt{{sink}}}}\rangle$ in a head set. The task plan is obtained by extracting the task vertices $v_{\texttt{TE}}.v_{\mathcal{T}}$ from each $v_{\texttt{TE}}$ along $\Pi_{v_{\texttt{TE}}^{\texttt{{src}}}v_{\texttt{TE}}^{\texttt{{sink}}}}$. The resulting task plan is an unvalidated schedule, which does not account for the motion feasibility along the task transitions.
In our formulation, task constraints are defined as $\mathcal{C}_{\mathcal{T}}=\mathcal{C}_{h}\cup\mathcal{C}_{f}$, where $\mathcal{C}_{f}=(v_{\mathcal{T}}^{\text{pre}},v_{\mathcal{T}}^{\text{post}})$ denotes frontier constraints and $\mathcal{C}_{h}=(v_{\mathcal{T}}^{\text{pre}},v_{\mathcal{T}}^{\text{post}})$ denotes history constraints. A frontier constraint specifies that $v_{\mathcal{T}}^{\text{pre}}$ must not be present in the frontier for $v_{\mathcal{T}}^{\text{post}}$ to be expanded, while a history constraint specifies that $v_{\mathcal{T}}^{\text{pre}}$ must not appear in the history for $v_{\mathcal{T}}^{\text{post}}$ to be expanded. Frontier constraints tie hyperarc expansion to the current state, whereas history constraints depend on the previous states.
This formulation naturally captures geometric constraints: for example, an object placement in the current state may block the expansion of another task-space element, but once the object is moved, the expansion becomes feasible. Moreover, the constraints are chainable, meaning a $v_{\mathcal{T}}^{\text{post}}$ can serve as a $v_{\mathcal{T}}^{\text{pre}}$ for another constraint. We enforce these constraints by restricting hyperarc expansion, preventing the formation of partial hyperarcs whenever a prohibited vertex exists in the head set of the currently forming hyperarc.
@@
IV-C Task Conflict Detection Layer
This section presents the second layer of the hierarchy, the task conflict detection layer, which validates the minimum feasibility of the unvalidated schedule by tracing transition configurations of robots and static objects throughout the schedule. An overview of this layer is provided in Algorithm 2 and Figure 3.
The unvalidated schedule assumes that all motions within the task plan are feasible, leaving the subsequent motion planning layer to confirm motion feasibility. However, relying solely on the motion planning layer to uncover all constraints (e.g., geometric constraints) becomes computationally expensive as the number of robots and objects increases.
IV-C1 Task Conflict
To mitigate this, we identify task conflicts by sampling the minimum configuration information required to evaluate task-level feasibility. This includes the sequence of object poses across the schedule and the corresponding robot grasp configurations computed via inverse kinematics, guided by the formation constraint $f_{c}$ in the task vertex $v_{\mathcal{T}}$. These samples specify the explicit configurations $q$ at each transition point, which are then traced through the schedule to detect collisions. Detected conflicts are represented as motion vertex–vertex conflicts, where each motion vertex encodes the configuration of robots and objects at a transition in the unvalidated schedule. For example, such conflicts may arise when two objects occupy colliding positions, or when a robot in a grasping configuration collides with an unrelated object due to obstruction. Such conflicts are validated throughtout the entire unvalidated schedule to collect all conflicts existing in the current task plan, to entirely capture the potential task constraints.
IV-C2 Task Constraint Feedback
When collisions involve objects that can be resampled (e.g., not anchored at start or goal poses), the conflict is resolved by generating alternative object poses within the planning area, which is handled directly in the task conflict detection layer. In contrast, if a manipulator collides with an obstructing object, the resolution requires replanning by propagating task constraints back to the task query phase. The identified constraints captures the entirety of the conflicts in the current unvalidated schedule and can be reused throught the task query phase afterwards.
However, task constraints may conflict when the order of task space vertices is reversed, leading to contradictions. To resolve this, additional stable poses must be introduced as move-out states, requiring new task space elements, since the minimal initial representation may be insufficient to satisfy all detected constraints during the task query process. In such cases, the task conflict detection layer identifies the object involved in the unsatisfiable constraints and expands the task space representation with an additional stable pose, thereby expanding the hypergraph to include the new transitions.
The iterative loop of querying, detecting task constraints, and expanding the task space continues until a task-level feasible plan is generated, which is then forwarded to the motion planning layer for further validation.
Algorithm 2 Task Conflict Detection
1:Unvalidated schedule $\Pi_{v_{\texttt{TE}}^{\texttt{{src}}}v_{\texttt{TE}}^{\texttt{{sink}}}}$, task-space hypergraph $\mathcal{H}_{\mathcal{T}}$
2:Task constraint set $\mathcal{C}_{\mathcal{T}}$
[…]
5:for $E_{\texttt{TE}}\in\Pi_{v_{\texttt{TE}}^{\texttt{{src}}}v_{\texttt{TE}}^{\texttt{{sink}}}}$
6:  $q\leftarrow\textsc{GetTransitionConfigs}(f,E_{\texttt{TE}}.E_{\mathcal{T}})$
7:  $C\leftarrow\textsc{CollisionCheck}(f,q)$
8:  if $C\neq\emptyset$
9:   $\mathcal{C}_{\mathcal{T}}\leftarrow\mathcal{C}_{\mathcal{T}}\cup C$
10:  $\textsc{ApplyAction}(f,E_{\texttt{TE}}.E_{\mathcal{T}}.\texttt{Tail})$
11:return $\mathcal{C}_{\mathcal{T}}$
Figure 3: Illustration of task conflict detection. In the left figure, $O_{1}$ and $O_{2}$ collide at frontier 3, creating a task constraint that blocks the expansion of hyperarcs with $O_{1}$ at their head. Expansion of such hyperarcs becomes possible only when $O_{2}$ is absent at the frontier, in accordance with the task constraints.
IV-D Motion Planning Layer
@@
This section describes the third layer of the hierarchy, the motion planning layer, which involves the construction of the motion representation and the query phases within the motion-level domain.
IV-D1 Representation Construction (Motion Hypergraph)
The motion hypergraph $\mathcal{H}_{\mathcal{M}}=(\mathcal{V}_{\mathcal{M}},\mathcal{E}_{\mathcal{M}})$ captures the motion details of $\mathcal{H}_{\mathcal{T}}$.
Each motion vertex $v_{\mathcal{M}}=\langle v_{\mathcal{T}},q\rangle\in\mathcal{V}_{\mathcal{M}}$ consists of a task-space vertex $v_{\mathcal{T}}$ and an explicit configuration $q$ sampled either by inverse kinematics or by the motion planner. This representation links configurations with the associated bodies, configuration spaces, and constraints defined in the task-space vertices. Each motion hyperarc $E_{\mathcal{M}}=\langle\texttt{Tail},\texttt{Head},E_{\mathcal{T}},\pi\rangle\in\mathcal{E}_{\mathcal{M}}$ is defined by its tail and head sets, the corresponding task hyperarc $E_{\mathcal{T}}$, and a configuration path $\pi$ that connects the motion vertices in $E_{\mathcal{M}}.\texttt{Tail}$ to those in $E_{\mathcal{M}}.\texttt{Head}$. As in DaSH [2], the motion hyperarcs can be categorized as a composition $E_{\mathcal{M}}^{\texttt{comp}}$, transition $E_{\mathcal{M}}^{\texttt{tran}}$, or move hyperarc $E_{\mathcal{M}}^{\texttt{move}}$. In our problem definition, $E_{\mathcal{M}}^{\texttt{tran}}$ denotes grasping or handover motions. For example, in a pick transition, the tail configurations include the object’s pose and the robot’s grasp configuration, each associated with a task-space vertex. The head configuration is the composite state of the robot grasping the object, where the grasp pose is constrained by the formation constraint $f_{c}$ in the task-space vertex $v_{\mathcal{T}}$. Any motion involving either a empty-handed robot or a robot holding an object is represented as $E_{\mathcal{M}}^{\texttt{move}}$.
Since the configurations at transition points (e.g., robot’s grasp pose) are already sampled and verified in the task conflict detection layer to provide the minimum information needed for task plan validation, these configurations directly specify the tail and head sets of $E_{\mathcal{M}}^{\texttt{tran}}$. This creates anchor points that the motion query must visit, bridging the $E_{\mathcal{M}}^{\texttt{move}}$ generated from different task space compositions. Roadmaps are then generated to serve as the basis for planning both $E_{\mathcal{M}}^{\texttt{tran}}$ and $E_{\mathcal{M}}^{\texttt{move}}$ during the subsequent motion query phase. In the Lazy-DaSH framework, these roadmaps are generated in an edge-unvalidated form following the principles of Lazy-PRM [11], with the feasibility of each motion deferred to validation during the subsequent query phase.
However, transition configurations may turn out to be invalid if they result in collisions with the environment, in which case resampling is performed to search for valid alternatives. For example, the transition configurations for a handover operation can be resampled by considering alternative mid-air poses of the object to avoid collisions with the environment. If the resampling attempts fail within the given budget, meaning no collision-free configuration exists to compose the $E_{\mathcal{M}}^{\texttt{tran}}$, the transition is considered infeasible. The planner then introduces task constraints that prohibit the corresponding task sequence, thereby preventing the expansion of the task hyperarc $E_{\mathcal{T}}$ and excluding the infeasible interaction during task plan replanning.
IV-D2 Query (Motion-extended Hypergraph)
Since the abstract guideline of the motion plan is defined by the task plan, the motion-extended hypergraph is basically a motion-detailed version of the task plan. The feasibility of the motions between these transition configurations is validated while tracing the task plan. The motion feasibility of task plan reflects the feasibility of $E_{\mathcal{M}}.\pi$ along the unvalidated schedule and is sequentially validated in a lazy manner by querying a lazy sampling-based motion planner such as Lazy-PRM [5]. The lazy motion validation approach minimizes unnecessary computational effort by deferring validation until it becomes necessary. If a motion query fails while tracing the unvalidated schedule, local roadmap improvements are attempted on the edges where coarse motion validation has succeeded within the given budget (e.g., timeout). If these improvements fail, additional vertices are sampled globally in the roadmap to search for an alternative global path. However, if motion planning still fails after a certain amount of effort (e.g., exceeding the timeout or the maximum number of roadmap improvement iterations), the process falls back to the task query phase, which backtracks to explore an alternative unvalidated schedule. Importantly, the invalid motion hyperarc is not permanently excluded from the motion hypergraph, allowing it to be revisited if it appears again under a different task plan.